OpenStudy (anonymous):

The volume of the solid obtained by rotating the region enclosed by y = x^2, \ y = 6 x, about the line x = 6 can be computed using the method of disks or washers via an integral

OpenStudy (luigi0210):

Do you know how to start?

OpenStudy (anonymous):

Yeah I know it starts from pi times the integral from 0 to 36 but i'm trying to figure out my two radiuses within the integral, all i know is that it has to relate to x=1/6y and x=sqrt(y)

zepdrix (zepdrix):

Ooo I wouldn't recommend doing that. I would just integrate in x. Have you learned the method of `shells`? It'll work really nicely. If you insist on integrating y I guess that's fine though.. -_-

OpenStudy (anonymous):

No I know how to use shells, it's just the problem requires me to use disks or washers

zepdrix (zepdrix):

ah :)

OpenStudy (anonymous):

Yeah so could you explain as to what my big Radius and little radius would be in this problem?

zepdrix (zepdrix):